# Federated Retrieval-Augmented Generation: A Systematic Mapping Study

### v1

@@ Abstract
Abhijit Chakrabortythanks: Corresponding author1  and Chahana Dahal2  and Vivek Gupta11footnotemark: 11 1Arizona State University   2Westminster University {achakr40,vgupt140}@asu.edu
@@ 1 Introduction
FL enables distributed model training across devices or institutions. Each client trains a model locally and only shares model updates (e.g., gradients), not raw data with each other. Common techniques like Federated Averaging (FedAvg) (McMahan et al., 2017) and secure aggregation (Bonawitz et al., 2017) help combine these updates safely. FL has been applied in healthcare, edge devices (e.g., smartphones), and finance, where data privacy is critical (Kairouz et al., 2021).
@@ 2 Methodology
We conducted a systematic literature search to map the research landscape at the intersection of federated systems and retrieval-augmented generation (RAG). Our search spanned top venues in natural language processing (e.g., ACL, EMNLP), machine learning (e.g., NeurIPS, ICLR), security and privacy (e.g., IEEE TIFS), as well as preprint servers like arXiv and relevant industry publications. We used combinations of keywords such as “federated learning”, “retrieval-augmented generation”, and “federated search”, and applied backward reference tracing to ensure inclusion of pivotal works cited in foundational papers (e.g., FedRetriever (Wang et al., 2024a), FLERAG (Mao et al., 2025)).
The search was restricted to works published between 2020 and 2025, reflecting the emergence and evolution of RAG in recent years. Out of an initial set of approximately 50 candidates, we identified 18 primary studies that explicitly addressed the integration of retrieval-based generation into federated or distributed settings. We excluded papers that focused solely on FL or RAG in isolation, without considering their combined architecture, training mechanisms, or deployment context.
Each selected work was systematically categorized along three dimensions, following Kitchenham’s guidelines for systematic mapping studies (Kitchenham et al., 2011): (a) Research Focus: The core technical challenge or system-level objective, such as privacy and security (e.g., secure retrieval, encrypted indexing), retrieval efficiency (e.g., resource selection and latency optimization), model/architecture integration (e.g., coupling RAG with FL pipelines), or domain-specific adaptation (e.g., tailoring systems for healthcare or QA). (b) Contribution Type: The nature of the technical contribution, such as a new model or framework, benchmark dataset, evaluation study, or conceptual survey. (c) Application Domain: The target deployment environment or intended use case, including clinical question answering, legal/financial reasoning, recommender systems, and multi-lingual enterprise settings.
@@ 3 Analysis of Recent Studies
Figure 2 categorizes Federated RAG architectures across retrieval strategies, generation mechanisms, and privacy/security enhancements. Retrieval strategies vary from simple naïve aggregation, suitable for small-scale scenarios with trusted participants, to selective query routing (e.g., RAGRoute (Guerraoui et al., 2025)), which reduces redundancy but adds complexity and potential accuracy risks. Secure retrieval using cryptographic protocols (e.g., FRAG (Zhao, 2024)) provides strong privacy protections but introduces computational overhead, making it ideal primarily for highly sensitive contexts.
Generation methods include centralized generation, which ensures consistency but risks privacy breaches and single points of failure, and client-specific generation, which offers superior personalization and data sovereignty at the expense of complexity in maintaining coherence across clients. Hybrid or ensemble strategies attempt to balance these extremes, combining central knowledge with localized adaptations, though increasing management complexity.
@@ 3.1 Research Focus
From 2023 through 2025, privacy mechanisms became system-level. C-FedRAG (Addison et al., 2024) executes retrieval and generation entirely inside trusted enclaves, securing clinical QA on MKP-QA. FRAG (Zhao, 2024) realises IND-CPA-secure $k$NN search via single-key HE while keeping latency practical.
@@
Table 1 summarises these trajectories across the 18 primary studies in our map. Half still prioritise privacy and security, reflecting the field’s FL heritage. Personalisation (e.g., GPT-FedRec, MKP-QA) and retrieval efficiency (e.g., RAGRoute, FedB4RAG) are rising, while model-level integration (FedE4RAG, UniMS-RAG) signals a shift toward deployable cross-silo systems.
Overall, research has progressed from algorithmic proofs of concept (2020–2022) to holistic architectures (2023–2025) that simultaneously safeguard data, route queries intelligently, and harmonise retrieval with generation, marking a decisive step toward scalable Federated RAG.
@@ 3.2 Contribution Type
Beyond system-level proposals, several contributions focus on datasets and benchmarking. FeB4RAG(Wang et al., 2024b) presents the first federated retrieval benchmark across 16 BEIR-derived domains, enabling rigorous evaluation of routing strategies and retrieval quality under resource constraints. MKP-QA(Shojaee et al., 2025) introduces a domain-specific benchmark for enterprise multilingual QA, designed to assess cross-silo document selection and context preservation. Additionally, RAGRoute (Guerraoui et al., 2025) offers a specialized evaluation testbed for its adaptive query planning mechanism, allowing performance measurement across silos under constrained budgets.
@@ 4 Research Activity and Publication Trends
As shown in Figure 3, publication volume increased significantly in 2024, with at least eight new studies emerging across venues in NLP (e.g., Findings of ACL), ML systems (e.g., MLSys, NeurIPS workshops), secure computing (e.g., IEEE TIFS), and industry white papers. This surge is not just numerical but also reflects broader shifts discussed in previous sections, specifically the diversification of research focus (Table 1) and emergence of novel models, benchmarks, and real-world applications (Table 2, Table 3). The steep rise in 2024 corresponds to two converging trends:
(i) The expanding use of LLMs in high-stakes domains such as healthcare and finance, where hallucination and outdated information are critical concerns.
(ii) Increasing regulatory and ethical scrutiny over centralized data processing, which incentivizes federated architectures.
Despite this momentum, the limited number of relevant studies observed in 2023 likely reflects the early-stage nature of Federated RAG research at that time, characterized by initial exploration of architectural concepts and limited practical deployments, preceding the subsequent significant increase in publication activity observed from 2024 onward, mostly appeared as preprints or workshop papers, with limited presence in archival venues. Notably, MIRAGE (Xiong et al., 2024) became one of the first benchmarks accepted to Findings of ACL 2024, signaling early consolidation around evaluation protocols. This aligns with trends seen in earlier paragraph on contribution type 3.2, where evaluation and analysis are gaining traction alongside system development.
Industry engagement is also increasing. C-FedRAG  (Addison et al., 2024) from NVIDIA proposes TEE-based confidential QA, while Adobe’s MKP-QA  (Shojaee et al., 2025) focuses on multilingual enterprise retrieval. Meanwhile, open-source initiatives like the Vector Institute’s FedRAG toolkit point to growing developer interest in practical deployment. Together, these developments reinforce the trends visualized in Figure 3, validating the observations made in the application domain and contribution type analyses. We expect Federated RAG to evolve into a distinct research track within NLP and AI by 2025–2026, underscoring the relevance of this systematic mapping study for guiding future work.
20202021202220232024202500555510101010# of Publications
@@ 4.1 Privacy & Security
Federated RAG must secure both retrieval and generation. C-FedRAG confines the entire pipeline to SGX enclaves, shielding queries and documents from untrusted servers. FRAG encrypts vector search with IND-CPA–secure homomorphic $k$NN, avoiding raw-index exposure, while FedE4RAG performs local augmentation at inference to minimise data movement. These techniques uphold regulations such as “right-to-be-forgotten” but add latency and can hamper cross-silo reasoning; future work should formalise threat models and quantify the privacy–utility frontier.
@@ 6 Challenges, Gaps and Trends
System Heterogeneity, Scalability, and Performance Trade-offs – Federated RAG systems must navigate heterogeneity in data distribution, storage capacity, and computational power across clients. Unified retrieval pipelines face challenges due to non-IID data and inconsistent resource availability. Naively querying all silos introduces latency and noise. To address this, RAGRoute employs lightweight classifiers for selective query routing, reducing source queries by up to 75% while maintaining answer quality. Similarly, the FeB4RAG benchmark highlights the importance of intelligent source selection and merging strategies in improving retrieval efficacy. Managing performance trade-offs, such as recall versus latency, is critical. Recent work like FRAG introduces encrypted caching and probabilistic gating to optimize resource use while maintaining privacy. Overall, these advances address architectural and efficiency concerns central to scalable federated RAG (RQ1).
@@ 7 Conclusion
Why this mapping study matters – Our study comprehensively maps the developing area of federated RAG to identify the technical gaps that hinder application in healthcare, banking, education, and cross-enterprise cooperation, where data sovereignty is non-negotiable but knowledge sharing is crucial. These problems require shared benchmarking testbeds, robust threat models, and lightweight on-device adapters to expand access to federated RAG to enable trustworthy AI in these environments. Federated RAG can grow from a promising niche to a foundation for secure, knowledge-rich language technology by bridging NLP, security, and distributed systems.
@@
Our findings are based on a targeted but relatively small number of primary studies. While this reflects the nascent state of the field, it also means that statistical generalizations or trend extrapolations should be interpreted with caution. For instance, the observed increase in 2024 publications may reflect a short-term spike due to LLM popularity rather than a long-term trajectory. Additionally, the evolving definitions and tooling of federated learning and RAG (e.g., hybrid decentralization models, model-serving architectures) suggest that future iterations of the field may adopt different terminology or structural assumptions. As such, the taxonomy and classifications presented here represent a snapshot of a fast-changing landscape.
@@
Selection bias remains a potential risk: our search protocol, while comprehensive, may have inadvertently excluded studies that use unconventional terminology or fall at the periphery of federated and retrieval-based learning. Furthermore, because many contributions in this domain come from industry, the absence of open-source implementations or full technical disclosures limits reproducibility and may skew the perceived landscape toward academic publications. Finally, while we have captured the most salient developments up to early 2025, the pace of innovation in LLM-based systems warrants periodic re-evaluation to maintain relevance and coverage.
Acknowledgment
We gratefully acknowledge the Complex Data Analysis and Reasoning Lab at Arizona State University for their resources and computational support.

### v2

@@ Abstract
Abhijit Chakrabortythanks: Corresponding author1  and Chahana Dahal2  and Vivek Gupta11footnotemark: 11 1Arizona State University   2University of Nevada, Las Vegas {achakr40,vgupt140}@asu.edu {chahana.dahal}@unlv.edu
@@ 1 Introduction
FL trains models across clients without sharing raw data, using methods like FedAvg (McMahan et al., 2017) and secure aggregation, widely applied in privacy-sensitive (Kairouz et al., 2021) domains.
@@
Federated RAG is distinctive because federated learning safeguards training-time privacy by restricting data to local silos, while retrieval-augmented generation grounds inference-time outputs in external evidence, reducing hallucinations. Their integration enables capabilities such as source attribution, local index maintenance, and dynamic document unlearning that neither paradigm achieves alone.
@@ 2 Methodology
We searched the best NLP, ML and security venues (2020-2025) using terms such as “federated learning”, “retrieval-augmented generation” and “federated search”, including backward references. From  50 papers, 18 met criteria. Following Kitchenham’s (Kitchenham et al., 2011) method, we coded each by (a) research focus, (b) contribution type, (c) application domain.
@@ 3 Analysis of Recent Studies
Retrieval ranges from naive aggregation to secure encrypted search (e.g., FRAG (Zhao, 2024)). Generation spans centralized, client-specific, or hybrid (e.g., GPT-FedRec (Zeng et al., 2024)). These choices balance scalability, personalization, and privacy (Figure 2).
@@ 3.1 Research Focus
From 2023 through 2025, privacy mechanisms became system-level. C-FedRAG (Addison et al., 2024) executes retrieval and generation entirely inside trusted enclaves, securing clinical QA on MKP-QA. FRAG (Zhao, 2024) realises IND-CPA-secure kkNN search via single-key HE while keeping latency practical.
@@
Table 1 summarizes these trajectories in the primary studies 18 on map. Half still prioritize privacy and security, reflecting the FL heritage of the field. Personalisation (e.g., GPT-FedRec, MKP-QA) and retrieval efficiency (e.g., RAGRoute, FedB4RAG) are rising, while model-level integration (FedE4RAG, UniMS-RAG) signals a shift toward deployable cross-silo systems.
Overall, research has progressed from algorithmic proofs of concept (2020–2022) to holistic architectures (2023–2025) that simultaneously safeguard data, route queries intelligently, and harmonise retrieval with generation, marking a decisive step toward scalable Federated RAG. To complement these qualitative descriptions, Table 5 in the Appendix synthesizes empirical metrics reported across representative systems, highlighting trade-offs between privacy, efficiency, and personalization.
@@ 3.2 Contribution Type
Beyond system-level proposals, several contributions focus on datasets and benchmarking. FeB4RAG(Wang et al., 2024b) presents the first federated retrieval benchmark across 16 BEIR-derived domains, enabling rigorous evaluation of routing strategies and retrieval quality under resource constraints. MKP-QA(Shojaee et al., 2025) introduces a domain-specific benchmark for multilingual enterprise QA, designed to assess the selection of cross-silo documents and the preservation of context. Additionally, RAGRoute (Guerraoui et al., 2025) offers a specialized evaluation testbed for its adaptive query planning mechanism, allowing performance measurement across silos under constrained budgets.
@@ 3.3 Application Domains and Use Cases
Beyond these domain categorizations, several privacy-critical scenarios further illustrate the practical relevance of Federated RAG. In healthcare, systems such as C-FedRAG and MIRAGE enable hospital QA across silos, allowing institutions to answer clinical questions from shared medical knowledge while ensuring that patient records never leave local storage. In financial services, approaches like RAGRoute support cross-division portfolio analysis, where sensitive compliance and investment documents remain within organizational boundaries yet can be queried collaboratively. Finally, in personal and consumer-facing settings, on-device assistants equipped with Federated RAG can retrieve from local files or personal notes without exposing them to cloud servers, offering a concrete pathway to privacy-preserving personal AI.
These scenarios underscore that Federated RAG extends beyond abstract architectural design to address urgent real-world concerns, where the dual imperatives of factual accuracy and data sovereignty are non-negotiable. By grounding retrieval and generation within privacy-preserving frameworks, such applications demonstrate the transformative potential of Federated RAG in bridging regulatory, ethical, and operational requirements.
3.4 Unified Objective Function and Architecture
While the taxonomy in Figure 2 organizes existing designs across retrieval, generation, and privacy axes, it is also useful to formalize how these components interact within a single optimization framework. We introduce a conceptual objective for Federated RAG see Figure 3 that captures joint training across clients:
$L_{\text{FedRAG}}(\theta)=\sum_{i=1}^{M}\tfrac{n_{i}}{N}\big{(}L_{\text{retrieve}}^{(i)}(\theta_{r})+L_{\text{generate}}^{(i)}(\theta_{r},\theta_{g})\big{)}$
where $M$ is the number of clients, $n_{i}$ is the number of samples on client $i$, NN is the total number of samples, $\theta_{r}$ are retriever parameters, and $\theta_{g}$ are generator parameters. This loss highlights that retrieval and generation can be jointly optimized in a federated setting, balancing local personalization with global consistency.
Figure 3: Conceptual architecture of Federated RAG, showing local query processing and index management at each client, coordinated with global model updates.
This abstraction complements the taxonomy: retrieval modules may use selective routing or encrypted kkNN, generation may be centralized, client-specific, or hybrid, and privacy layers such as secure enclaves or differential privacy wrap the pipeline. Together, the unified loss and architecture diagram illustrate how system design choices can be framed within a coherent optimization perspective.
@@ 4 Research Activity and Publication Trends
Figure 4 shows a sharp rise in 2024, driven by high-stakes LLM use (healthcare/finance) and regulatory pressure. Industry (e.g., NVIDIA’s C-FedRAG (Addison et al., 2024), Adobe’s MKP-QA (Shojaee et al., 2025)) and open-source toolkits signal consolidation around benchmarks like MIRAGE (Xiong et al., 2024) and FeB4RAG.
2020202120222023202420250551010# of Publications
@@
Alongside this growth, the field is also beginning to consolidate around shared metrics and benchmarks, which provide clearer evaluation protocols and practical entry points for researchers and practitioners. Notable examples include the MIRAGE clinical QA benchmark (Xiong et al., 2024), the FeB4RAG federated BEIR benchmark (Wang et al., 2024b), the RAGAS evaluation toolkit for hallucination and consistency (Es et al., 2024), and the open-source FedRAG toolkit (Vector Institute). Together, these resources mark an inflection point: Federated RAG research is evolving from isolated system proposals to a field with standardized evaluation pipelines. A comprehensive summary of these toolkits and datasets is provided in Appendix 5.
@@ 4.1 Privacy & Security
Federated RAG must secure both retrieval and generation. C-FedRAG confines the entire pipeline to SGX enclaves, shielding queries and documents from untrusted servers. FRAG encrypts vector search with IND-CPA–secure homomorphic kkNN, avoiding raw-index exposure, while FedE4RAG performs local augmentation at inference to minimise data movement. These techniques uphold regulations such as “right-to-be-forgotten” but add latency and can hamper cross-silo reasoning; future work should formalise threat models and quantify the privacy–utility frontier.
@@ 6 Challenges, Gaps and Trends
Design Patterns and Practical Guidelines
Beyond high-level challenges, recent studies also surface recurring design patterns that can guide implementers in practice:
  • Retrieval. Systems employ either selective query routing (e.g., RAGRoute) to reduce redundant cross-silo traffic, or encrypted kkNN search (e.g., FRAG) to ensure confidentiality at the cost of higher latency.
  • Generation. Approaches range from centralized generation, which improves consistency but raises privacy concerns, to fully client-side generation, which preserves data sovereignty but increases coordination complexity. Hybrid designs (e.g., GPT-FedRec) combine global knowledge with local personalization.
  • Personalization. Adapter tuning and federated embedding alignment (e.g., FedE4RAG) allow lightweight personalization without overwhelming communication budgets, making them attractive for heterogeneous client settings.
  • Trade-offs. Every design choice involves balancing privacy, latency, and accuracy. For example, secure enclaves (C-FedRAG) provide strong protection but add overhead; selective routing reduces cost but may risk coverage. These trade-offs highlight the importance of aligning system design with domain-specific regulatory and operational requirements.
Taken together, these patterns provide actionable guidance: implementers can view Federated RAG architectures not as ad hoc designs but as configurable combinations of retrieval, generation, and personalization strategies, each with well-understood implications for scalability and compliance.
System Heterogeneity, Scalability, and Performance Trade-offs – Heterogeneous clients introduce latency, uneven resource use, and noisy retrieval results. To address this, recent work explores selective query routing (e.g., RAGRoute), benchmark-driven source selection strategies FeB4RAG, and encrypted caching with probabilistic gating FRAG. Together, these methods illustrate how different design choices balance recall, privacy, and latency, highlighting the inherent trade-offs in federated RAG. Overall, such advances directly target the architectural and efficiency concerns that must be solved for scalable deployment (RQ1).
@@ 7 Conclusion
Why this mapping study matters –Federated Retrieval-Augmented Generation has rapidly progressed from early conceptual proposals to a diverse body of system designs, benchmarks, and evaluation resources. Our systematic mapping of 18 primary studies charts this trajectory, identifying common architectural patterns, key challenges, and emerging opportunities across domains such as healthcare, finance, and enterprise knowledge management.
In moving beyond description, this paper evolves from a static mapping to a field-shaping foundation: it combines structured analysis with empirical metrics, introduces theoretical perspectives through a unified objective function, and distills implementation guidance into actionable design patterns. Together, these contributions provide both researchers and practitioners with a clearer roadmap for advancing Federated RAG toward scalable, trustworthy, and domain-ready deployment.
@@
Our results stem from a small set of studies, so generalizations should be cautious; the 2024 surge may be a short-term spike. Evolving definitions and tools mean our taxonomy is a snapshot of a rapidly changing landscape.
@@
Despite a comprehensive search, some studies using unconventional terms or industry-only contributions may have been missed, limiting reproducibility. Given the rapid pace of LLM innovation, periodic re-evaluation is needed to keep this mapping relevant.
Appendix A Extended Resources and Comparative Synthesis
A.1 Benchmarks and Toolkits
To complement the taxonomy and trends discussed in the main paper, we provide expanded descriptions of benchmarks and toolkits that now support Federated RAG evaluation and deployment:
  • C-FedRAG (Addison et al., 2024): A confidential QA system executed entirely in Trusted Execution Environments (SGX), demonstrating strong privacy guarantees in clinical domains.
  • RAGRoute (Guerraoui et al., 2025): A query routing framework that reduces redundant cross-silo queries by up to 75% while maintaining accuracy.
[…]
  • GPT-FedRec (Zeng et al., 2024): A hybrid recommendation model combining collaborative filtering with RAG, achieving +5–7% personalization hit-rate.
  • FedE4RAG (Mao et al., 2025): A modular federated dense retriever framework with homomorphic encryption and distillation for end-to-end privacy.
  • UniMS-RAG (Wang et al., 2024a): A unified multi-source RAG system integrating client-side retrieval with a shared generator for adaptive latency control.
  • MIRAGE (Xiong et al., 2024): A clinical QA benchmark designed for federated hospital data, formalizing evaluation in healthcare contexts.
  • MKP-QA (Shojaee et al., 2025): A benchmark for enterprise multilingual QA across product silos, supporting federated evaluation in industry.
  • FeB4RAG (Wang et al., 2024b): A federated benchmark spanning 16 BEIR-derived domains to study routing strategies and retrieval efficiency.
  • RAGAS (Es et al., 2024): An evaluation toolkit introducing hallucination, factuality, and robustness metrics tailored to federated RAG pipelines.
  • FedRAG (Addison et al., 2024): A distributed EHR QA system that introduces hierarchical retrieval across hospitals while preserving patient privacy.
  • RAG (Wang et al., 2024a): A medical large language model integrated with RAG, demonstrating gains over FL-only baselines.
  • DP-FedKGE (Peng et al., 2021): A framework for differentially private federated knowledge graph embeddings, enabling reasoning with entity-level privacy.
  • PPDA (Jeon et al., 2021): A lightweight decentralized aggregation protocol that ensures privacy without relying on DP or HE.
  • DeTrustFL (Xu et al., 2022): A cryptographic trust consensus mechanism preventing disaggregation attacks in federated setups.
  • FedX (Han et al., 2022): An unsupervised federated representation learning system using cross-client knowledge distillation.
  • FL–LLM surveys (Wu et al., 2025): A survey of federated LLM fine-tuning methods, with explicit discussion of RAG integration.
  • FRAD (Zhao) (Zhao, 2024): A general-purpose federated architecture demonstrating cryptographic retrieval under multi-domain constraints.
Together, these resources provide the beginnings of a shared evaluation layer, enabling more consistent comparisons across systems and domains.
A.2 Full Comparative Synthesis of Surveyed Studies
Table 5 extends the synthesis presented in Section 3.1 by aligning each of the 18 Federated RAG studies (2020–2025) with their reported benefits and outcomes. This consolidated view highlights where empirical improvements have been quantified, where cryptographic and privacy guarantees dominate, and where new benchmarks or toolkits have been introduced.
Appendix B Comparative Synthesis of Federated Studies
In addition to the qualitative taxonomy presented in Section3.1, we provide a consolidated comparative synthesis of all 18 Federated RAG studies surveyed between 2020–2025 (Table5). This extended table aligns each system with its reported benefit and empirical or conceptual outcome, thereby offering a more granular view of how diverse contributions map to performance, privacy, efficiency, and evaluation. For instance, system-level advances such as C-FedRAG and RAGRoute quantify measurable improvements in accuracy and communication efficiency, while cryptographic frameworks like FRAG demonstrate the latency trade-offs inherent in strong privacy guarantees. Benchmarking and evaluation initiatives, including FeB4RAG, MIRAGE, and RAGAS, highlight the field’s growing emphasis on standardized protocols, whereas systems like FedE4RAG and UniMS-RAG focus on architectural integration. Early foundational efforts (e.g., DP-FedKGE, PPDA, DeTrustFL) continue to shape privacy-preserving mechanisms, while recent domain-specific benchmarks (e.g., MKP-QA) expand application horizons. Taken together, this synthesis complements the main text by grounding the survey’s classifications in concrete evidence and by illustrating the breadth of technical strategies underpinning Federated RAG research.
System + Reported Benefit Outcome / Contribution
C-FedRAG  (Addison et al., 2024)
 Clinical QA in trusted enclaves +12.7% QA accuracy (59.8→72.5) using SGX
RAGRoute (Guerraoui et al., 2025)
 Adaptive query routing across silos –75% redundant queries with 72% accuracy
FRAG (Zhao, 2024)
[…]
GPT-FedRec (Zeng et al., 2024)
 Personalized recommendation +5–7% hit rate improvement
FedE4RAG (Mao et al., 2025)
 Federated dense retrievers + HE distillation End-to-end privacy-preserving retrievers
UniMS-RAG (Wang et al., 2024a)
 Multi-source retrieval + shared generator Unified retrieval-control, adaptive latency
MIRAGE (Xiong et al., 2024)
 Clinical QA benchmark First federated benchmark for clinical QA
MKP-QA (Shojaee et al., 2025)
 Enterprise multilingual QA benchmark Cross-silo multilingual evaluation
FeB4RAG (Wang et al., 2024b)
 Federated BEIR benchmark Evaluation across 16 BEIR-derived domains
RAGAS (Es et al., 2024)
 Evaluation toolkit for hallucination/consistency Metrics for hallucination, evidence quality, robustness
FedRAG (Addison et al., 2024)
 Distributed EHR QA Hierarchical hospital retrieval pipeline
RAG (Wang et al., 2024a)
 Medical LLM + RAG integration Outperforms FL-only baselines on clinical QA
DP-FedKGE (Peng et al., 2021)
 Differentially private embeddings Entity-level DP protection for KG embeddings
PPDA (Jeon et al., 2021)
 Lightweight decentralized aggregation Privacy without DP/HE overhead
DeTrustFL (Xu et al., 2022)
 Cryptographic trust consensus Prevents disaggregation attacks in FL
FedX (Han et al., 2022)
 Cross-client knowledge distillation Unsupervised federated representation learning
FL–LLM surveys (Wu et al., 2025)
 Survey of FL-LLM fine-tuning Overview, references to RAG methods
FRAD (Zhao) (Zhao, 2024)
 General multi-domain architecture Cryptographic constraints, scalable retrieval
Table 5: Comparative synthesis of 18 Federated RAG studies (2020–2025). Each system (bold, first line) is paired with its reported benefit (indented line) and the corresponding outcome or contribution.